Engineers are working on a problem observed with Atlantis’ External Tank (ET-127), which will likely require the removal and replacement of a valve.Mold impressions of the damaged GH2 two inch disconnect valve were taken overnight, with meetings taking place on Friday to discuss a forward plan. However, should R&R be required, work is expected to be completed within processing milestone timelines.
**The most comprehensive collection of Shuttle, Ares, Orion and ISS related presentations and mission documentation, plus expansive daily processing documentation and updates are available to download on L2 **
**Click here for sample of L2 menu and content**
**STS-124 L2 Sub Section (over 3400mb) – Continuing through post flight and post mission IFA/Mission reviews. 13 Post Flight IFA Review Presentations now available.
**STS-125/LON-400 L2 Sub Section Special already over 1750mb in size – special section Now Live, with over 40 superb mission presentations. STS-125 Flight Plan (newly created) and Flight Data Files now uploaded. Mission Overview in CGI videos building up – starting with payload overviews. STS-400 Flight Plan. STS-400 EVA and mission overview Presentations (14). Daily extensive processing latest.
**STS-124 Post Flight – STS-119 Processing Latest** — **STS-124 News Articles**
**LIVE updates: Endeavour Dual LON and STS-126 Processing**
**LIVE updates on Atlantis STS-125 Processing Flow** — **STS-125 News Articles**
ET-127 Valve Issue:
ET-127 arrived at the Kennedy Space Center (KSC) earlier in the week, following a superb production effort at the Michoud Assembly Facility (MAF). Upon arriving in the Vehicle Assembly Building (VAB), the tank was lifted into High Bay 4’s checkout cell for pre-mating checks.
During overnight checks, the GH2 two inch disconnect valve was found to be damaged. A report of the issue was already on the desk of shuttle managers by the morning.
‘External Tank: ET-127 (VAB HB-4 Checkout Cell): Inspection and rework continue. The Ground Umbilical Carrier Plate (GUCP) 7 inch and 3/8inch quick disconnects (QD) are installed,’ noted processing information on Friday.
‘Aft Hardpoint Closeout preps are complete with primer in cure. Vent Valve Actuation Flex Hose installation is work complete.
‘The GH2 2 inch disconnect is under Engineering evaluation. Mold impressions were taken and read last night and will be passed to Boeing and Orbiter this morning. Assessment is in work for the potential R&R impact.’
Follow-up notes confirm that any requirement to replace is not expected to delay the processing milestones, which pick up at the end of the month with the mating of the ET to STS-125’s twin Solid Rocket Boosters (SRBs). R&R work will be carried out by Boeing engineers, if required.
The GH2 disconnect valve is located at the bottom end of the tank on the LH2 umbilical plate. Its function is to provide gaseous hydrogen from the orbiter’s Main Propulsion System (MPS) to the ET hydrogen tank. This allows for the pressurization of the tank once the propulsion system is up and running.
The gaseous hydrogen is supplied to the feed system from a low pressure hydrogen pump – fed from hot LH2 emitting off the engine coolant system. The operating pressure is about 35-40 PSIG.
There is also a similar system for LO2 – though that is supplied with oxygen gas from a dedicated heat exchanger on the SSMEs – which is specifically designed for that purpose.
UPDATE: R&R of the valve has been approved.
Atlantis/STS-125 Processing Latest:
Meanwhile, processing of Atlantis remains issue free, as she inches ever closer to rollover to the VAB next month. Inspections are being carried out in the OPF (Orbiter Processing Facility), with a large amount of ‘sensitive cleaning’ – to remove dust and particles from flight hardware.
‘Bent pin inspections were performed yesterday with no anomalies found. Hubble Space Telescope mid-body highly sensitive cleaning continues through next week. Port TSA tool installation was completed yesterday less safety wire and conformal coating,’ added Friday processing information.
**New L2 Trailer Video, with examples of L2 content – TRAILER NOW AVAILABLE – CLICK IMAGE**. **Includes FIVE Stunning high quality 2hr, 355-400mb Camcorder and HUD videos, with crew audio – from payload bay closure – through re-entry with an astronaut held camcorder video – to post landing – several more videos showing landing from 90,000 ft also available and HUD videos from STA landings.
‘RCC (Reinforced Carbon Carbon) thermography inspections were completed yesterday. Rudder/speedbrake (RSB) thermal barrier R&R was completed yesterday. RSB closeout inspections today will be followed by manual closure, shaft installation, and vertical closeout.
‘Sky Genie line inspections were completed yesterday. Orbiter and OMS closeouts for flight have begun and continue into next week. Side hatch cycle and latch gap checks will be performed today.
‘SSME carrier panel installations, flipper door closures, and FRCS (Forward Reaction Control System) electrical access door installations will be worked today. Wing leading edge sensor battery installation and checkout will occur today.
‘Weekend work: On Saturday, HST (Hubble Space Telescope) highly sensitive cleaning, ECS monthly maintenance, and TPS work. HST cleaning may roll into Sunday based on progress.’
Endeavour/STS-400/STS-126 Processing Latest:
Over in OPF-2, parallel processing is continuing with Endeavour’s flow towards supporting STS-125 via Launch On Need STS-400. Highlights include the completion of SSME installation activities and work on the RSB tab/clips – a result of the liberation event during the latter stages of STS-124’s flight.
‘SSME-mounted and dome heat shields are installed for flight. The nose landing gear was raised yesterday for more mylar pulls. Orbiter Ku-band DTV interface testing was completed yesterday,’ added processing information.
‘Rudder/speedbrake spring tabs were removed yesterday; installation today. Sky genie descent line inspection was pulled in on Wednesday and completed. RCC panel 10R thermography is scheduled for today.
‘APU 3 (Auxiliary Power Unit) electrical functional testing picks up today. Fuel cell water re-servicing is expected to resume today. External airlock forward hatch functional is scheduled for today. Weekend work: LH (Left Hand) ET door tile bonding and splash on the stinger door tiles on Saturday.’
A pico satellite was also installed into Endeavour on Thursday – which is now undergoing interface verification testing today.
Discovery/STS-119 Processing Latest:
In OPF 3, Discovery is also enjoying a smooth flow, with work continuing on the installation of a special experiment to her Thermal Protection System.
Originally scheduled to be conducted on Endeavour’s STS-126 mission later this year, the Boundary Layer Transition DTO (Detailed Test Objective) was moved to STS-119, due to processing constraints on Endeavour ahead of her unique LON (Launch On Need) support role for STS-125.
The Boundary Layer Transition DTO will involve the use of ‘integrated TPS and MADS (Modular Auxiliary Data System) instrumentation (designed) to fly with a fixed protuberance tile with a height of 0.25” to initiate BLT at around Mach 15,’ noted the documentation.
To help mitigate some of the risk associated with this DTO, a total of three TPS tiles downstream of the protuberance will be changed to eliminate the risk of overheating in these areas. In addition to this mitigation effort, thirteen temperature sensors will be installed in surrounding TPS tiles.
‘A total of thirteen thermocouples will be installed in TPS tiles: eleven will be in the affected wing area downstream of the transitional cone of influence, one in the protuberance tile, and one as a reference upstream of the protuberance.’
Of the thirteen sensors, ten will be selected for temperature recording during flight. This recording will take place through ‘the MADS instrumentation system via wiring to a new thermal reference junction (TRJ) in the wing and from there (through) wiring (most existing) to the mid-body MADS shelf,’ added the baseline presentation.
In all, 22 tiles and 20 gap fillers will be affected during TPS modifications for this DTO.
‘MPS 12-inch flow liner inspections continue through Monday. The -Z and -Y star tracker door removals were completed yesterday. The expended drag chute pyros was removed yesterday,’ added other notes of interest on processing information.
‘Forward hatch seal R&R was completed on 2nd shift last night. Hatch will be installed today. The CDR/PLT inertial reel seat modification was completed yesterday. Shuttle Orbiter Repackaged Galley (SORG) functional testing was completed yesterday. Weekend work: TPS waterproofing on Sunday.’